You are planning to buy a bond and are considering 2 options. Your one requirement is that you need to buy the lower price bond. Your desired nominal yield after tax is 4%, convertible semi-annually.

The details of the 2 bonds under consideration are as follows:

  • A corporate bond with a 10,000 par value, a 4 year term, and a coupon rate of 6% payable semi-annually. The tax rate applicable on corporate bond coupon payments is 30%. The bond will be held to maturity, at which time a capital gain or loss is realized. The capital gains tax rate is 0%.
  • A municipal bond with a 10,000 par value, a 4 year term, and a coupon rate of 4% payable semi-annually. The tax rate on municipal bond coupon payments is zero. The bond will be held to maturity, at which time a capital gain or loss is realized. The capital gains tax rate is 0%.

          Find the price of both bonds. Which one will you purchase?
Answer #1

For half year interest payments, the periods got double and interest rates half.

Formula: coupon * tax effect * PVIFA(Yield, periods) + Par value * PVIF(Yield, periods)
Price of Corporate bond= 10000*6%*0.7/2 *PVIFA(2%,8) + 10000 * PVIF(2%,8)
210*7.325 + 10000*0.853 = 10068.25
Price of Municipal bond= 10000*2% *PVIFA(2%,8) + 10000 * PVIF(2%,8)
200*7.325 + 10000*0.853 = 9995
We will purchase Municipal bonds being more safer returns and present lower price.